OpenStudy (anonymous):

Graph the exponential model y=3(6)^x. Which point lies on the graph? A. (-1, -18) B. (1, 18) C. (2, 9) D. (-9, 2)

OpenStudy (anonymous):

if \(x=1\) then \(3\times 6^1=3\times 6=18\) so for sure \((1,18)\) is on the graph

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw